Erratum for “On the étale homotopy type of Morel-Voevodsky spaces”

Alexander Schmidt

Juli 2004

1. As Ofer Gabber pointed out to me, Remark 4.5 is incorrect. Fortunately, this does not affect anything else in the paper.

2. There where some discussions on the validity of [MV] 2.1.13 (=Proposition 2.1 of this paper) because of the rather sketchy proof there. Meanwhile, Fabien Morel has sent me a complete proof (and Rick Jardine further simplified the argument).